# SemiOccam: A Robust Semi-Supervised Image Recognition Network Using Sparse Labels

### v2

@@ 3.3 Pseudo-labeling Mechanism
4:  Initialize $\{\pi_{l},\bm{\mu}_{l},\Sigma_{l}\}$ from cluster results
@@
6:     EM algorithm: Initial iterative via (4-9)
@@
8:  Construct $\mathcal{D}_{p}$ using (11-13) with $\tau$ and $\alpha$
@@
11:     EM algorithm: Final iterative via (4-9)
@@ 4.1 Standard SGMM Performance Analysis
Figure 3: Relationship between number of labels/PCA dimensions/Gaussian components and performance on the MNIST dataset. The first figure shows a steady improvement in accuracy with an increase in labels. The second figure shows the trend of SGMM accuracy with PCA dimensions, while the right figure shows the trend of accuracy with the number of Gaussian submodels.
@@ 4.2 More Powerful Feature Extractors
To further analyze the performance of the SGMM, we experimented on the CIFAR-10 dataset using different feature extractors: PCA, Deep Neural Networks (ResNet101, DLA169), and ViT-Base/Large backbone DINO pre-trained models. Figure 4 shows that with DNNs and ViTs, increasing the number of Gaussian components leads to a decrease in classification accuracy, which contrasts with the performance trend observed when using PCA.
@@
Through experimental results and visualization analysis, we found that DNN and ViT can better capture the distribution characteristics of data compared to the traditional PCA dimensionality reduction method, thereby improving the classification performance of SGMM.
4.3 Experiment Setup
Datasets.
We report results on the Cifar-10/100 and CleanSTL-10 datasets. CIFAR-10 contains 10 classes with 6,000 32x32 color images per class, totaling 60,000 images, of which 50,000 are for training and 10,000 for testing. CIFAR-100 is similar to CIFAR-10 but contains 100 classes with 600 images per class, of which 500 are for training and 100 for testing. STL-10 is an image classification dataset designed for self-supervised and semi-supervised learning, containing 10 classes with 5,000 labeled training samples and 100,000 unlabeled images, each at a resolution of 96×96 pixels. We created CleanSTL-10 by carefully removing duplicates and overlaps between training and test sets, as detailed in Appendix A. After this deduplication, CleanSTL-10 contains 90,455 training images and 8,000 testing images, ensuring more reliable and fair evaluation.
Implementation details.
We use the DINO pre-trained model based on the ViT-large backbone to extract feature representations. On the Cifar-10/100 datasets, the values of the Gaussian components are set to 10/100 respectively, and the PCA dimension is set to 60. In the STL-10 dataset, since its training set contains more than 10 categories, we set the number of Gaussian components to 15 to obtain better robustness, the PCA dimension is set to 45, and the convergence threshold (Tol) is set to 1. The PCA dimension is set to use accuracy as an evaluation metric to evaluate the classification performance of different methods. All experiments are performed on a Tesla T4 GPU with 15GB memory.
4.4 Comparison with State-of-the-art Methods
Results on CIFAR-10.
We show the benchmark results on CIFAR-10 in Table 1. Our method achieves competitive performance compared with state-of-the-art methods, including FixMatch, UDA, and ReMixMatch. It can be clearly seen that our method consistently outperforms other methods under all settings.
Table 1: Benchmark results on CIFAR-10.
Dash (2021) 9.29 $\pm$ 3.28 5.16 $\pm$ 0.23 4.36 $\pm$ 0.11
MPL (2021) 6.62 $\pm$ 0.91 5.76 $\pm$ 0.24 4.36 $\pm$ 0.11
FlexMatch (2021) 4.97 $\pm$ 0.06 4.98 $\pm$ 0.09 4.19 $\pm$ 0.01
CoMatch (2021) 6.51 $\pm$ 1.18 5.35 $\pm$ 0.14 4.27 $\pm$ 0.12
SimMatch (2022) 5.38 $\pm$ 0.01 5.36 $\pm$ 0.08 4.41 $\pm$ 0.07
AdaMatch (2022) 5.09 $\pm$ 0.21 5.13 $\pm$ 0.05 4.36 $\pm$ 0.05
FreeMatch (2023) 4.90 $\pm$ 0.12 4.88 $\pm$ 0.09 4.16 $\pm$ 0.06
SoftMatch (2023) 5.11 $\pm$ 0.14 4.96 $\pm$ 0.09 4.27 $\pm$ 0.05
SequenceMatch (2024a) 4.80 $\pm$ 0.01 4.75 $\pm$ 0.05 4.15 $\pm$ 0.01
EPASS (2024b) 5.55 $\pm$ 0.21 5.31 $\pm$ 0.13 4.23 $\pm$ 0.05
SemiOccam(Ours) 3.51 $\pm$ 0.12 3.47 $\pm$ 0.17 3.45 $\pm$ 0.16
Results on CleanSTL-10.
Table 2: Benchmark results on CleanSTL-10.
Dash (2021) 42.00 $\pm$ 4.94 10.50 $\pm$ 1.37 6.30 $\pm$ 0.49
MPL (2021) 35.97 $\pm$ 4.14 9.90 $\pm$ 0.96 6.66 $\pm$ 0.00
FlexMatch (2021) 29.12 $\pm$ 5.04 9.85 $\pm$ 1.35 6.08 $\pm$ 0.34
CoMatch (2021) 13.74 $\pm$ 4.20 7.63 $\pm$ 0.94 5.71 $\pm$ 0.08
SimMatch (2022) 16.98 $\pm$ 4.24 8.27 $\pm$ 0.40 5.74 $\pm$ 0.31
AdaMatch (2022) 19.95 $\pm$ 5.17 8.59 $\pm$ 0.43 6.01 $\pm$ 0.02
FreeMatch (2023) 28.50 $\pm$ 5.41 9.29 $\pm$ 1.24 5.81 $\pm$ 0.32
SoftMatch (2023) 22.23 $\pm$ 3.82 9.18 $\pm$ 0.63 5.79 $\pm$ 0.15
SequenceMatch (2024a) 15.45 $\pm$ 1.40 12.78 $\pm$ 0.76 5.56 $\pm$ 0.35
EPASS (2024b) 9.15 $\pm$ 3.25 6.27 $\pm$ 0.03 5.40 $\pm$ 0.12
SemiOccam(Ours) 4.57 $\pm$ 0.24 4.43 $\pm$ 0.08 4.15 $\pm$ 0.07
Results on CIFAR-100.
Table 3: Benchmark results on CIFAR-100.
Dash (2021) 44.82 $\pm$ 0.96 27.15 $\pm$ 0.22 21.88 $\pm$ 0.07
MPL (2021) 46.26 $\pm$ 1.84 27.71 $\pm$ 0.19 21.74 $\pm$ 0.09
FlexMatch (2021) 39.94 $\pm$ 1.62 26.49 $\pm$ 0.20 21.90 $\pm$ 0.15
CoMatch (2021) 53.41 $\pm$ 2.36 29.78 $\pm$ 0.11 22.11 $\pm$ 0.22
SimMatch (2022) 39.32 $\pm$ 0.72 26.21 $\pm$ 0.37 21.50 $\pm$ 0.11
AdaMatch (2022) 38.08 $\pm$ 1.35 26.66 $\pm$ 0.33 21.99 $\pm$ 0.15
FreeMatch (2023) 39.52 $\pm$ 0.01 26.22 $\pm$ 0.08 21.81 $\pm$ 0.17
SoftMatch (2023) 37.60 $\pm$ 0.24 26.39 $\pm$ 0.38 21.86 $\pm$ 0.16
SequenceMatch (2024a) 37.86 $\pm$ 1.07 25.99 $\pm$ 0.22 20.10 $\pm$ 0.04
EPASS (2024b) 38.88 $\pm$ 0.24 25.68 $\pm$ 0.33 21.32 $\pm$ 0.14
SemiOccam(Ours) 26.59 $\pm$ 1.02 22.19 $\pm$ 0.81 21.21 $\pm$ 0.26
4.5 Training Cost Analysis
Our proposed method utilizes a pre-trained ViT as a feature extractor. After this feature extraction step is completed, the core SGMM classifier training process can be completed in less than ten minutes on a free Tesla T4 GPU 15GB on the Google Colab platform, demonstrating high efficiency and state-of-the-art performance. This minute-level training time gives our method a significant competitive advantage in resource-constrained practical deployments and rapid iteration research scenarios.
4.6 Ablation Study
Comparing Semi-supervised Learning Methods.
In order to test the effectiveness of the SGMM component, we compared it with the original classification head of ViT. We trained on three datasets, error rates shown in Table 4, under scenarios with extremely few labels, the performance of SGMM is significantly superior to that of the Softmax layer.
Effect of Pseudo-labeling.
We conducted an ablation study to evaluate the impact of pseudo-labeling on the performance of SemiOccam. We compared the performance of SemiOccam with and without pseudo-labeling on three datasets, error rates shown in Table 4.
Table 4: Ablation study with three datasets.
Dataset CIFAR-10 CIFAR-100 CleanSTL-10
# Label 40 250 4000 400 2500 10000 40 250 1000
5 Conclusions
We introduce high-performance feature extractors into the classical yet often underappreciated MoE classifier for image recognition tasks. Our architecture is simple and training is highly efficient, making it particularly well-suited for scenarios with limited labeled data. The proposed unified modeling strategy performs surprisingly well when trained with very few labeled samples and large-scale unlabeled data. In our study, we also identify a long-overlooked data leakage issue in the STL-10 dataset and release the deduplicated CleanSTL-10 dataset to ensure fair and reliable evaluation in semi-supervised learning. Extensive experiments show that SemiOccam achieves new performance breakthroughs on benchmarks such as CleanSTL-10 and CIFAR-10, maintaining over 96% classification accuracy even with only 4 labeled samples per class in under ten minutes on a single Tesla T4 GPU, compared to the hundreds of GPU hours required by existing methods.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62076078 and in part by the Chinese Association for Artificial Intelligence (CAAI)-Huawei MindSpore Open Fund under Grant CAAIXSJLJJ-2020-033A.
APPENDIX
Appendix A STL-10 Data Clean
Duplicate samples in the STL-10 dataset can potentially bias model evaluation. To address this, we deduplicated the dataset. Specifically, we used a script to compute image hashes for STL-10 dataset, then removed 7,545 samples from the train set that were found to be duplicates of samples in the test set. We release the deduplicated CleanSTL-10 dataset to facilitate fair and reliable research. The pseudo-code for this deduplication process is provided in Algorithm A.1.
Algorithm A.1 Deduplication of Dataset
@@ Appendix A STL-10 Data Clean
We have compiled a table of the indices for training and test samples with duplicate hash values, which you can verify in our project repository.
Figure B.1: Heatmap comparison. From left to right: raw image, DLA169, DINO. Note the small penguin on the right in the third row.
Appendix B Comparing Heatmaps of DNN and ViT
We compare the heatmaps of DNN and ViT on several randomly selected images from the internet, each containing two animals simultaneously. The heatmaps are generated using the Grad-CAM method, which visualizes the importance of each pixel in the image for the classification result. As shown in Figure B.1, the heatmaps of DNN and ViT differ significantly. While DNN focuses primarily on one animal, ViT attends to both animals in the image, demonstrating its ability to capture global context more effectively. This comparison highlights the differences in how these models prioritize regions of the image during classification.
Appendix C PCA Explained Variance
In our experiments, the hyperparameter of PCA dimensions for different datasets is determined by the PCA explained variance, which represents the percentage of variance in the data that can be explained by the feature vectors after PCA dimensionality reduction. We typically choose dimensions that can explain over 60% of the variance.
Figure C.1: PCA explained variance. The percentage of variance explained by each PCA dimension on the CIFAR-10 dataset.
Appendix D ViT Classification Head for Semi-Supervised Learning
The following is the mathematical implementation of the original ViT classification head for semi-supervised learning, which has been used in our ablation study in section 4.6.
In section 3.1, we define the feature vector $\mathbf{f}_{i}\in\mathbb{R}^{d}$. The standard ViT classification head includes a linear transformation $\mathbf{s}_{i}=\mathbf{W}_{h}\mathbf{f}_{i}+\mathbf{b}_{h}$, where $\mathbf{W}_{h}\in\mathbb{R}^{K\times d}$ and $\mathbf{b}_{h}\in\mathbb{R}^{K}$ are learnable parameters, and $K$ is the number of categories; followed by softmax:
@@ Appendix D ViT Classification Head for Semi-Supervised Learning
For labeled samples $\mathcal{F}_{l}=\{(\mathbf{f}_{i},c_{i})\}$, we optimize the standard cross-entropy loss:
$\mathcal{L}_{\text{sup}}=-\frac{1}{|\mathcal{F}_{l}|}\sum_{(\mathbf{f}_{i},c_{i})\in\mathcal{F}_{l}}\log P(c=c_{i}|\mathbf{f}_{i})$
Unlabeled data $\mathcal{F}_{u}=\{\mathbf{f}_{j}\}$ adopts a confidence-based pseudo-labeling strategy, where the pseudo-label is $\hat{c}_{j}=\arg\max_{k}P(c=k|\mathbf{f}_{j})$, and the confidence score is $\xi_{j}=\max_{k}P(c=k|\mathbf{f}_{j})$. The unsupervised loss is calculated as follows:
$\mathcal{L}_{\text{unsup}}=-\frac{1}{|\mathcal{F}_{u}|}\sum_{\mathbf{f}_{j}\in\mathcal{F}_{u}}\mathbb{I}(\xi_{j}>\tau)\log P(c=\hat{c}_{j}|\mathbf{f}_{j})$
The total loss combines the supervised and unsupervised objectives, with the weight $\lambda(t)$ changing over time, where $T_{\text{ramp}}$ corresponds to the training period:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sup}}+\underbrace{\lambda_{\text{max}}\cdot\min\left(1,\frac{t}{T_{\text{ramp}}}\right)}_{\lambda(t)}\mathcal{L}_{\text{unsup}}$

### v3

@@ 3.3 Pseudo-labeling Mechanism
4:  Initialize $\{\pi_{l},\boldsymbol{\mu}_{l},\boldsymbol{\Sigma_{l}}\}$ from cluster results
@@
6:     EM algorithm: Initial iterative via (6-11)
@@
8:  Construct $\mathcal{D}_{p}$ using (13-15) with $\tau$ and $\alpha$
@@
11:     EM algorithm: Final iterative via (6-11)
@@ 4.1 Standard SGMM Performance Analysis
Figure 4: Relationship between number of labels/PCA dimensions/number of gaussian components and performance on the MNIST dataset. The first figure shows a steady improvement in accuracy with an increase in labels. The second figure shows the trend of SGMM accuracy with PCA dimensions, while the right figure shows the trend of accuracy with the number of gaussian submodels.
@@ 4.2 More Powerful Feature Extractors
To further analyze the performance of the SGMM, we experimented on the CIFAR-10 dataset using different feature extractors: PCA, DNNs (ResNet101, DLA169), and ViT-Base/Large backbone DINO pre-trained models. Figure 5 shows that with DNNs and ViTs, increasing the number of gaussian components leads to a decrease in classification accuracy, which contrasts with the performance trend observed when using PCA.
@@
We compare the heatmaps of DNN and ViT on several randomly selected images from the internet, each containing two animals simultaneously. The heatmaps are generated using the Grad-CAM method, which visualizes the importance of each pixel in the image for the classification result. As shown in Figure 8, the heatmaps of DNN and ViT differ significantly. While DNN focuses primarily on one animal, ViT attends to both animals in the image, demonstrating its ability to capture global context more effectively. This comparison highlights the differences in how these models prioritize regions of the image during classification. Through experimental results and visualization analysis, we found that DNN and ViT can better capture the distribution characteristics of data compared to the traditional PCA dimensionality reduction method, thereby improving the classification performance of SGMM.
Figure 8: Visual comparison of feature attention (left) and PCA variance analysis (right).
4.3 STL-10 Data Clean
Duplicate samples in the STL-10 dataset can potentially bias model evaluation. To address this, we deduplicated the dataset. Specifically, we used a script to compute image hashes for STL-10 dataset, then removed 7,545 samples from the train set that were found to be duplicates of samples in the test set. We release the deduplicated CleanSTL-10 dataset to facilitate fair and reliable research. The pseudo-code for this deduplication process is provided in Algorithm 2.
Algorithm 2 Deduplication of Dataset
@@ 4.3 STL-10 Data Clean
Table 1: Example pairs of training and test samples sharing the same hash value.
Train Index Test Index Hash
5008 2075 e777922f14273159776a089a6f44d2bbb03f23a2b618f0914ef7ca442e770396
5025 4149 d91a18f5a643aaf3569387641a6f40a92ec82a32a27a532625414dbaf48b4d25
5032 6188 eadb8e0cd40606bb3e6c58f710abb4d9769e43b43724ef764e7a74a4d179fff4
5041 4228 5c1dee15a65d67508008125e0f1085382084a860e491c10f0f6c15d7c7d06a36
5045 6029 f4197466cd0d97da2a2b079bc7458d7af8d1c89a7178d30d14668364037d165a
5062 4980 82a4e1da419ea969f5bc07f9664e8974dee8c8161334d0ad9abc89f9545e8f41
5071 7166 b445c22c7f8505e1ac09d6efa5b756daac42da5ae70acd8d672b9c164c4c8bbc
5072 7091 9fb660ef9603189ca18787bcf60966bc8cb07f6ad121eeb3602cb543a5123830
5073 2803 7e3753d43ea8769a2818d947feea2beb84cf5e3cb090d5b31c820e6f6f4f0d8a
5092 6820 c360dd72cb9f9056302cafbc2f5ccfe9de66fddac0b8ed4d91ce93996dd98503
5108 5761 03a02e10956c65799a5f5d082d61cbaf1de414773c7c025da8bfbc5553c6fdf0
5121 2143 8e98699fcf6ec4e0b8765866248e7702c2f26f9fe2d1f469eeb0f51f54bb5afb
5129 1287 363360e93d8997c1138d1e4a0fde65b7b08befd496b450df72989aac5e5b4edb
5139 3879 b2fa77caa967437f1a85c6071b2345d420da69780036d44f761591ad6d213ff6
5143 4701 29b8233f6d9a24cfbbdd83f23c838d59b1a113ada232c0991bd63e1fa97c533e
5165 6844 46f4be3edd44d770964fb27f6777e77f5a93d1623921c9d630f9454908f9e2ac
5169 4665 ee864c70c6d01868bcde6eb61261c33057f39efbe9197a7590aed629a4215bae
5172 461 aa81e6c7a9a3c597b331abd352271f91eb4f1129b41cffd164b35371c41a8217
5188 2893 0da766b7fe175d6cb8328437019e4fa2d87f4878cdb09e9b2d001fe6374a100e
5191 4131 b909207b8b5822d2a2197ab4563fc32def9b9d04934285928a0ad6f08d1de9f4
5197 6931 013a41f75629ce371fbbec3f9a845aa5ec3e615819a4e4ae67b3dbe0d70ece0d
5214 991 9a3201824495feabca12e1ef2339c410ff1f30c412ff35d85743bcd16f6bcabc
5226 196 7309f100d4dc411520175d905e2cdebdd3ace6422ec0310efb61f7a93052402b
@@
We present an example table of training and test sample indices with duplicate hash values in Table 1. The full list can be found in our project repository.
4.4 Experiment Setup
Datasets.
We report results on the Cifar-10/100 and CleanSTL-10 datasets. CIFAR-10 contains 10 classes with 6,000 32x32 color images per class, totaling 60,000 images, of which 50,000 are for training and 10,000 for testing. CIFAR-100 is similar to CIFAR-10 but contains 100 classes with 600 images per class, of which 500 are for training and 100 for testing. STL-10 is an image classification dataset designed for self-supervised and semi-supervised learning, containing 10 classes with 5,000 labeled training samples and 100,000 unlabeled images, each at a resolution of 96×96 pixels. We created CleanSTL-10 by carefully removing duplicates and overlaps between training and test sets, as detailed in Section 4.3. After this deduplication, CleanSTL-10 contains 90,455 training images and 8,000 testing images, ensuring more reliable and fair evaluation.
Implementation details.
We use the DINO pre-trained model with a ViT-Large backbone to extract feature representations. For dimensionality reduction, PCA is applied with dimensions selected based on explained variance, typically preserving over 60% of the total variance (see Fig. 8). On the CIFAR-10 and CIFAR-100 datasets, we set the number of gaussian components to 10 and 100, respectively, and the PCA dimension to 60. For CleanSTL-10, which contains more than 10 categories in its training set, we set the number of gaussian components to 15 for improved robustness, the PCA dimension to 45, and the convergence threshold (Tol) to $10^{1}$. Classification performance is evaluated using accuracy. All experiments are conducted on a single NVIDIA Tesla T4 accelerator with 15 GB of memory.
4.5 Comparison with State-of-the-art Methods
Results on CIFAR-10.
We show the benchmark results on CIFAR-10 in Table 2. Our method achieves competitive performance compared with state-of-the-art methods, including FlexMatch [39], SequenceMatch [19] and EPASS [20]. It can be clearly seen that our method consistently outperforms other methods under all settings.
Table 2: Benchmark results on CIFAR-10.
Dash (2021) 9.29 $\pm$ 3.28 5.16 $\pm$ 0.23 4.36 $\pm$ 0.11
MPL (2021) 6.62 $\pm$ 0.91 5.76 $\pm$ 0.24 4.36 $\pm$ 0.11
FlexMatch (2021) 4.97 $\pm$ 0.06 4.98 $\pm$ 0.09 4.19 $\pm$ 0.01
CoMatch (2021) 6.51 $\pm$ 1.18 5.35 $\pm$ 0.14 4.27 $\pm$ 0.12
SimMatch (2022) 5.38 $\pm$ 0.01 5.36 $\pm$ 0.08 4.41 $\pm$ 0.07
AdaMatch (2022) 5.09 $\pm$ 0.21 5.13 $\pm$ 0.05 4.36 $\pm$ 0.05
FreeMatch (2023b) 4.90 $\pm$ 0.12 4.88 $\pm$ 0.09 4.16 $\pm$ 0.06
SoftMatch (2023) 5.11 $\pm$ 0.14 4.96 $\pm$ 0.09 4.27 $\pm$ 0.05
SequenceMatch (2024a) 4.80 $\pm$ 0.01 4.75 $\pm$ 0.05 4.15 $\pm$ 0.01
EPASS (2024b) 5.55 $\pm$ 0.21 5.31 $\pm$ 0.13 4.23 $\pm$ 0.05
SemiOccam (Ours) 3.51 $\pm$ 0.12 3.47 $\pm$ 0.17 3.45 $\pm$ 0.16
Results on CleanSTL-10.
Table 3: Benchmark results on CleanSTL-10.
Dash (2021) 42.00 $\pm$ 4.94 10.50 $\pm$ 1.37 6.30 $\pm$ 0.49
MPL (2021) 35.97 $\pm$ 4.14 9.90 $\pm$ 0.96 6.66 $\pm$ 0.00
FlexMatch (2021) 29.12 $\pm$ 5.04 9.85 $\pm$ 1.35 6.08 $\pm$ 0.34
CoMatch (2021) 13.74 $\pm$ 4.20 7.63 $\pm$ 0.94 5.71 $\pm$ 0.08
SimMatch (2022) 16.98 $\pm$ 4.24 8.27 $\pm$ 0.40 5.74 $\pm$ 0.31
AdaMatch (2022) 19.95 $\pm$ 5.17 8.59 $\pm$ 0.43 6.01 $\pm$ 0.02
FreeMatch (2023b) 28.50 $\pm$ 5.41 9.29 $\pm$ 1.24 5.81 $\pm$ 0.32
SoftMatch (2023) 22.23 $\pm$ 3.82 9.18 $\pm$ 0.63 5.79 $\pm$ 0.15
SequenceMatch (2024a) 15.45 $\pm$ 1.40 12.78 $\pm$ 0.76 5.56 $\pm$ 0.35
EPASS (2024b) 9.15 $\pm$ 3.25 6.27 $\pm$ 0.03 5.40 $\pm$ 0.12
SemiOccam (Ours) 4.57 $\pm$ 0.24 4.43 $\pm$ 0.08 4.15 $\pm$ 0.07
Results on CIFAR-100.
Table 4: Benchmark results on CIFAR-100.
Dash (2021) 44.82 $\pm$ 0.96 27.15 $\pm$ 0.22 21.88 $\pm$ 0.07
MPL (2021) 46.26 $\pm$ 1.84 27.71 $\pm$ 0.19 21.74 $\pm$ 0.09
FlexMatch (2021) 39.94 $\pm$ 1.62 26.49 $\pm$ 0.20 21.90 $\pm$ 0.15
CoMatch (2021) 53.41 $\pm$ 2.36 29.78 $\pm$ 0.11 22.11 $\pm$ 0.22
SimMatch (2022) 39.32 $\pm$ 0.72 26.21 $\pm$ 0.37 21.50 $\pm$ 0.11
AdaMatch (2022) 38.08 $\pm$ 1.35 26.66 $\pm$ 0.33 21.99 $\pm$ 0.15
FreeMatch (2023b) 39.52 $\pm$ 0.01 26.22 $\pm$ 0.08 21.81 $\pm$ 0.17
SoftMatch (2023) 37.60 $\pm$ 0.24 26.39 $\pm$ 0.38 21.86 $\pm$ 0.16
SequenceMatch (2024a) 37.86 $\pm$ 1.07 25.99 $\pm$ 0.22 20.10 $\pm$ 0.04
EPASS (2024b) 38.88 $\pm$ 0.24 25.68 $\pm$ 0.33 21.32 $\pm$ 0.14
SemiOccam (Ours) 26.59 $\pm$ 1.02 22.19 $\pm$ 0.81 21.21 $\pm$ 0.26
4.6 Ablation Study
We conduct comprehensive ablation experiments to evaluate the effectiveness of different components in our SemiOccam framework. Table 5 summarizes the classification error rates across three datasets under various settings.
4.6.1 Effectiveness of SGMM vs. Softmax Classifier
To evaluate the benefit of using the SGMM component, we replace it with the standard ViT classification head based on a softmax layer and train both models under limited-label settings. The results in Table 5 show that SemiOccam significantly outperforms the softmax-based head, especially in extremely low-label regimes.
ViT Classification Head for Semi-Supervised Learning.
We define the original ViT head as follows: Given the feature vector $\mathbf{f}_{i}\in\mathbb{R}^{d}$ extracted from the backbone, a linear projection is applied as $\mathbf{s}_{i}=\mathbf{W}_{h}\mathbf{f}_{i}+\mathbf{b}_{h}$, followed by a softmax function:
@@ 4.6 Ablation Study
We adopt a confidence-based pseudo-labeling scheme to construct the unsupervised loss. The overall training objective combines both supervised and unsupervised loss terms:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{sup}}+\lambda(t)\cdot\mathcal{L}_{\text{unsup}},\quad\lambda(t)=\lambda_{\text{max}}\cdot\min\left(1,\frac{t}{T_{\text{ramp}}}\right)$
4.6.2 Impact of Pseudo-Labeling
We further examine the effect of pseudo-labeling by disabling it in training. The row labeled “w/o P-L” in Table 5 shows that removing pseudo-labeling leads to degraded performance, highlighting its importance for leveraging unlabeled data effectively.
Table 5: Ablation results on CIFAR-10/100 and CleanSTL-10.
Dataset CIFAR-10 CIFAR-100 CleanSTL-10
# Labels 40 250 4000 400 2500 10000 40 250 1000
4.7 Training Cost Analysis
Our proposed method utilizes a pre-trained ViT as a feature extractor. After this feature extraction step is completed, our core classifier training process can typically be completed in less than ten minutes on a single free Tesla T4 accelerator on the Google Colab platform, demonstrating high efficiency and state-of-the-art performance. This minute-level training time gives our method a significant competitive advantage in resource-constrained practical deployments and rapid iteration research scenarios. For detailed training time performance, please refer to Figure 9.
Figure 9: Our Core Training Time Analysis. This surface plot illustrates the training time (in seconds) of the SGMM classifier under varying numbers of GMM and PCA components, assuming a 60,000-sample training dataset.
5 Conclusions
We introduce high-performance feature extractors into the classical yet often underappreciated MoE classifier for image recognition tasks. Our architecture is simple and training is highly efficient, making it particularly well-suited for scenarios with limited labeled data. The proposed unified modeling strategy performs surprisingly well when trained with very few labeled samples and large-scale unlabeled data. In our study, we also identify a long-overlooked data leakage issue in the STL-10 dataset and release the deduplicated CleanSTL-10 dataset to ensure fair and reliable evaluation in semi-supervised learning. Extensive experiments show that SemiOccam achieves new performance breakthroughs on benchmarks such as CleanSTL-10 and CIFAR-10, maintaining over 96% classification accuracy even with only 4 labeled samples per class in under ten minutes on a single Tesla T4 accelerator, compared to the hundreds of GPU hours required by existing methods.
Acknowledgments
The work is supported by the National Natural Science Foundation of China No. 62076078, Fundamental Research Funds for the Central Universities No.3072024LJ0403, and the CAAI-Huawei MindSpore Open Fund No. CAAIXSJLJJ-2020-033A.